OpenStudy (anonymous):

simplify: x/2x+5=-5/x

OpenStudy (anonymous):

\[\frac{ x }{ 2x+5 } = \frac{ -5 }{ x }\] \[x ^{2} = -5 (2x+5)\] \[x^{2} = -10x -25\] \[x^{2} +10x +25 =0 \] \[(x+5)^{2} = 0\] you will stop at this step since we are only asked to simplify not solve x= -5 , -5

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

URW :)